OpenStudy (anonymous):

Graph each system of inequalities. Name the coordinates of the vertices of the feasible region. Find the maximum and the minimum values of the given function.

OpenStudy (anonymous):

\[5\ge y \ge-3 \] \[4x+y \le5\] \[-2x+y \le5\] \[f(x,y)=4x-3y\]